Use the following information (assume the tax rate is 35 percent):2014 2015 Sales Depreciation Cost of goods sold Other expenses Interest Cash Accounts receivable Short-term notes payable Long-term debt Net fixed assets Accounts payable Inventory Dividends $ 9,635$ 10,209 1,306 3,240 714 783 5,383 6,307 926 15,440 17,850 36,255 37,437 4,365 9,85010,118 1,231 1,305 2,876 819 705 4,289 5,619 974 4,676 1,136

For 2015, calculate the cash flow from assets, cash flow to creditors, and cash flow to stockholders.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Σ AutoSum ー E ゴWrap Text ta copy ▼ в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting, as Table w styles. ▼ ㆆ ▼ Sort &Find & 2 ClearFe Select Edting Format Painter Clipboard CG172 172 173 174 175 176 ncome statement SALES COGS OTHER EXPENSES DEPRECIATION EBIT INTEREST EBT TAX NET INCOME 10209 3240 714 1306 949 783 4166 1458.10 707.90 LESS LESS LESS 178 179 180 181 182 183 184 185 186 187 188 189 190 1 LESS OCF = EBIT + DEPRECIATION-TAXES OCF 4949 + 1306-1458, 10 = 4796.90 Cashflow from assets OCF NET CAPITAL SPENDING-NET CHANGE IN WORKING CAPITAL 4796.90-2488-2361- 52.10 | | OCF CASH FLOW STATW ARM, MSI,confidence INDEXCURRENCYLOAN CROSS OVER HEDGING MONEY MARKET GM HPR AM TAX GROWTH SUST PREF AND DIV POLICYI 06:51 21-01-2019 Cl 2273971 2 2 --LE--Home nert Page Layout Formulas Data Review View dd-Ins Cut Copy Σ AutoSum ー E ゴWrap Text в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard CI203 CB Alignment Number Cells Edting CD CE CF CG CH CI CJ CK CL CM 185 186 187 188 189 190 191 192 193 194 195 196 197 198 199 200 201 202 203 1 | | OCF CASH FLOW STATW OCF EBIT DEPRECIATION TAXES OCF = 4949 + 1306-1458.10 = 4796.90 Cashflow from assets = OCF NET CAPITAL SPENDING -NET CHANGE IN WORKING CAPITAL 4796.90 2488 2361- 52.10 NET CAPITAL SPENDING = ENDING NET FIXED ASSETS-BEGINNING FIXED ASSETS + DEPRECIATION 37437 -36255 1306 2488 NET CHANGE IN WORKING CAPITAL-CHANGE IN CURRENT ASSETS-CHANGE IN CURRENT LAI BILTIES NET CHANGE IN WORKING CAPITAL 2050 (-311) 2361 (5383+6307+10118)-(4289+5619+9850) (4365-4676) 2050 311 CHANGE IN CURRENT ASSETS = CHANGE IN CURRENT LIABILITIES = cashflow to creditors interest paid-net new borrowings including notes payable (783-(17850-15440)-(926-974)) -1579 ARM, MSI.confidence INDEX CURRENCYLOANCROSS OVER HEDGING MONEY MARKETGM HPR AM TAXGROWTH SUSTPREF AND DIV POLICY 06:51 21-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 C Paste Conditional Format CeInsert Delete Format Formatting as Table Styles2 Clear Sort &Find & Format Painter Clipboard CI217 CB Alignment Number Cells Edting CD CE CF CG CH CI CJ CK CL CM 199 200 201 202 203 204 205 206 207 208 209 210 211 212 213 214 215 216 217 1 | | OCF CASH FLOW STATW cashflow to creditors interest paid-net new borrowings including notes payable (783-(17850-15440)-(926-974)) -1579 cash flow from assets -cash flow to creditors +cashflow to stockholders [-52.1--1579+cashflow to stock holders] cashflow to stock holders- 1579 -52.1- 1526.9 ARM, MSI.confidence INDEX CURRENCYLOANCROSS OVER HEDGING MONEY MARKETGM HPR AM TAXGROWTH SUSTPREF AND DIV POLICY 06:51 21-01-2019